AIRSPEED It takes a private airplane 8.75 hours to make the
2,100-mile flight from Atlanta to Los Angeles and 5 hours to make the
return trip. Assuming that the wind blows at a constant rate from Los
Angeles to Atlanta, find the airspeed of the plane and the wind rate.

Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

It takes a private airplane 8.75 hours to make the 2,100-mile flight from Atlanta to Los Angeles and 5 hours to make the return trip.
Assuming that the wind blows at a constant rate from Los
Angeles to Atlanta, find the airspeed of the plane and the wind rate.
------------------
ATL to LA DATA:
time = 8.75 hrs ; distance = 2100 ; rate = d/t = 2100/8.75 = 240 mph
-----
LA to ATL DATA:
time = 5 hrs ; distance = 2100 ; rate = d/t = 2100/5 = 420 mph
-------
Equation:
With-wind:: p + w = 420 mph
Against ::: p - w = 240 mph
------------------------------
Add and solve for "p"::
2p = 660
plane = 330 mph (speed in still air]
-----
Subtract and solve for "w"::
2w = 180
w = 90 mph (wind rate)
-------------------------
Cheers,
Stan H.
--------------

Answer by Alan3354(69443)

It takes a private airplane 8.75 hours to make the
2,100-mile flight from Atlanta to Los Angeles and 5 hours to make the
return trip. Assuming that the wind blows at a constant rate from Los
Angeles to Atlanta, find the airspeed of the plane and the wind rate.
-----------------------
Find the 2 groundspeeds, upwind and downwind using
r = d/t
You get mi/hr if you divide the miles by the hours.
---
The plane's airspeed is the average of the 2 groundspeeds.
The windspeed is the difference between airspeed and groundspeed.